74 inch

All Systems Operational

Jean Bernardo IT Standby - Wednesday, June 3rd, 2020 at 17:45 SAST

No work was done today.

Brendt Christian Friday, May 29th, 2020 at 15:44 SAST

Put SpupNic instrument on vacuum

Brendt Christian Wednesday, May 27th, 2020 at 10:06 SAST

Did morning and evening rounds

Hilton Kamfer Monday, May 18th, 2020 at 17:47 SAST

Did morning and afternoon rounds

Hilton Kamfer Monday, May 18th, 2020 at 17:47 SAST

Did morning and afternoon rounds and filled the instrument with LN2.

Hilton Kamfer Sunday, May 17th, 2020 at 16:57 SAST

Did morning and afternoon rounds and filled the instrument with LN2.

Hilton Kamfer Saturday, May 16th, 2020 at 16:08 SAST

Did morning and afternoon rounds and filled the instrument with LN2.

Hilton Kamfer Friday, May 15th, 2020 at 11:44 SAST

Did morning and afternoon rounds and filled the instrument with LN2.

Hilton Kamfer Thursday, May 14th, 2020 at 10:54 SAST

Did morning and afternoon rounds and filled the instrument with LN2.

Brendt Christian Tuesday, May 12th, 2020 at 11:03 SAST

Did morning and afternoon rounds, made sure telescope is remotely operational and filled the instrument with LN2.

Brendt Christian Monday, May 11th, 2020 at 14:57 SAST

Did morning, evening rounds and filled SpupNic with LN2

Brendt Christian Sunday, May 10th, 2020 at 14:11 SAST

Did rounds in the morning, put Spupnic on vacuum and did afternoon rounds

Brendt Christian Saturday, May 9th, 2020 at 14:21 SAST

Did rounds

Brendt Christian Friday, May 8th, 2020 at 12:13 SAST

Did monthly maintenance service on dome and telescope. Checked if there is enough CO2 for mirror cleaning. We will need LN2, a trip for Monday 11/05/2020 has been arranged. One small dewer is filled which will be used after vacuuming the instrument.

Brendt Christian Thursday, May 7th, 2020 at 13:40 SAST

Did morning Rounds

Brendt Christian Wednesday, May 6th, 2020 at 11:17 SAST

Did morning rounds. Did testing, rotated the dome, opened and closed the shutter and tested, tested the mirror cover and tested the telescope moving it in RA and DEC.

Jean Bernardo IT Standby - Tuesday, May 5th, 2020 at 17:36 SAST

No work was done today.

Jean Bernardo IT Standby - Monday, May 4th, 2020 at 17:51 SAST

No work was done today.